# Unexpected fault activation in underground gas storage. Part II: Definition of safe operational bandwidths

### v1

@@ 3.1 Effect on rock and caprock
The interactions between gas, brine, and rock can lead to significant changes in the reservoir and fault properties, which, in turn, can greatly impact the performance, efficiency, and safety of the storage process. The impact of gas injection on the mechanical response of sedimentary rocks remains uncertain to date. Existing research predominantly focuses on CO2 injection (Kim and Makhnenko, 2022), with limited investigation into N2 injection (Fuchs et al., 2019). There is some ongoing development regarding H2 injection into sedimentary rocks. The latest comprehensive review on H2 injection and withdrawal has been compiled by Corina et al. (2022). From the available literature on CO2-fluid-rock interactions, e.g. Rohmer et al. (2016) and Peter et al. (2022), two contrasting viewpoints emerge: CO2 can either have a negligible influence on the rock mechanical parameters, or degrade them because of mineral dissolution and modifications in pore size distribution. A stable condition is often observed when the mineral composition and pore structure remain unchanged during CO2 injection (Rimmelé et al., 2010; Bolourinejad and Herber, 2015), with the main elastic properties largely unaffected by the CO2 presence (Mikhaltsevitch et al., 2014; Hu et al., 2016). On the other hand, in cases where a weakening effect is observed after CO2 injection, the mechanical strength of the rock decreases, leading to increased permeability and reduced rock stiffness. The dissolution of carbonate minerals, particularly calcite, reduces the cohesion of the rock matrix and weakens grain-to-grain contacts (Kim and Makhnenko, 2022). This effect has been observed in experiments on sedimentary rocks like sandstones and carbonates. Generally, carbonate-rich rocks are more susceptible to mechanical weakening due to CO2-fluid-rock interactions.
@@
While several experimental evidences support these findings, uncertainties still persist in the results. For instance, CO2-acidified brine injection can precipitate minerals, potentially increasing the strength and stiffness of the rock depending on the specific material and its mineral composition. CO2 can react with certain minerals like calcite, inducing mineral precipitation and cementation, thereby enhancing the mechanical integrity of the rock (Espinoza et al., 2018). Sandstones, for example, may experience both increase and decrease in the deformation moduli (Marbler et al., 2012). The variability in results regarding the effects on rock stiffness and strength highlights the need for additional experimental data to address existing uncertainties. Regarding the geochemical effects of CO2 injection in the Upper Rotliegend formation, Bolourinejad and Herber (2015) reported an increase in permeability ranging from 10% to 30%. This suggests changes in porosity and deformation moduli, with potential deformation moduli variation of $\pm$30%.
@@
Regarding N2 and H2 injection, there is a limited investigation. N2 is primarily used for determining rock porosity or as a control gas (Fuchs et al., 2019). Therefore, the exposure of rocks to N2 is expected to have negligible effects. In the case of H2 injection, quantitative data is lacking, but qualitative rock degradation is expected. Hydrogen injection in UHS porous reservoirs may trigger the dissolution of carbonates and sulfates, as well as grain crushing and local compaction, altering the reservoir porosity and mechanical and flow properties (Al-Yaseri et al., 2023). The response will be time-dependent, with pH equilibration to higher values potentially leading to further compaction and stiffening of the reservoir. However, this effect is likely to stabilize after a few cycles. Consequently, the poroelastic response of the reservoir will heavily depend on the chemical environment in the early stages of operations and become more stable and predictable in later stages.
3.2 Effect on faults
@@
From a goechemical point of view, gas injection may also affect faults, which represent the key factors responsible for fluid migration and containment. Gas injection can affect fault permeability, reactivation potential, and overall stability, which could lead to unintended fluid pathways or seismic activity. Concerning the CO2, during the storage phase over long periods (on the order of thousands of years), fluid-rock interactions can lead to changes in fault rock fabric and mineralogy. This may result in fault frictional properties and cohesion alterations due to dissolution and cementation processes (Collettini et al., 2008). However, laboratory friction tests comparing fault rocks exposed to CO2 for several hours to samples taken from natural analogs exposed to CO2 over millions of years, show limited friction reduction due to CO2 exposure (Samuelson and Spiers, 2012). Moreover, no specific laboratory studies show how CO2-enhanced dissolution and re-precipitation processes may impact on fault cohesion. Based on the literature, the chemical effects on fault mechanical properties after CO2 injection are considered as negligible.
@@
4 Model set-up
We use the mathematical and numerical framework developed, discussed and tested in Part I of this work (Franceschini et al., 2024). The modeling approach consists of one-way coupled Finite Element (FE) hydro-poromechanical simulations in fractured geological media. For all the details, the reader is referred to Franceschini et al. (2024).
Refer to caption Refer to caption
Figure 3: On the left: Plain view of the conceptual model. On the right: vertical sections of the conceptual model along the trace A-A and B-B.
@@ 4.1 Conceptual reference model
The domain geometry conceptually reproduces the compartmentalization commonly observed in the UGS fields of the Rotliegend formation and already sketched by Franceschini et al. (2024). The reservoir compartments are approximately located at the center of a 30$\times$30$\times$5 km domain (Figure 3). The reservoir is 200-m thick, 2-km deep, and composed by two adjacent 2$\times$2 km blocks. The two reservoir compartments are separated by a fault (F3) and confined by two sets of orthogonal faults (F1-F2 along the y-axis, F4-F5 along the x-axis). Depending on the sealing properties of fault F3, the two compartments can have a partial hydraulic connection. The faults extend from 3-km to 1.6-km depth and terminate in the Zeichestein salt formation sealing the reservoir on top. Faults F1 and F2 have a dip angle of ±$\pm 85^{\circ}$, while faults F4 and F5 are vertical. The dip angle of fault F3 can vary from $-65^{\circ}$ to $+65^{\circ}$. Block 2 can be moved up and down, with a maximum offset of 200 m with respect to the Block 1.
@@
The 3D domain is discretized by 8-node hexahedra, with 253,165 nodes and 236,208 elements. A finer discretization is used between 1800 and 2200 m depth. Here, the reservoir element charateristic size is 100$\times$100$\times$20 m. The fault system embedded in the 3D grid is discretized by 5,215 zero-thickness contact elements. Standard boundary conditions with zero displacements on the outer and bottom surfaces are prescribed, whereas the top surface, representing the ground, has a stress-free condition. Figure 4 shows a view of the discretized domain with the embedded fault system.
Figure 4: Axonometric view of the 3D computational grid used for the geomechanical simulation and the embedded 2D grid for the fault system.
@@
The setting of the reference scenario is established on the basis of the hydro-geomechanical properties of the Rotliegend formation. The fault F3 is kept vertical, with no offset between the reservoir blocks. Overburden, caprock, reservoir, and underburden are assumed to behave elastically, with the properties reported in Table 1. Faults are governed by Coulomb’s failure criterion with cohesion $c=2$ MPa and static friction angle $\varphi_{s}=30^{\circ}$. A linear slip-weakening is also possible with a reduction to the dynamic friction angle $\varphi_{d}$ at a sliding equal to $d_{c}$ (Franceschini et al., 2024). As to the undisturbed stress regime, we assume the vertical effective stress, $\sigma_{v}$, to be a principal component. The other two horizontal principal components of the effective stress tensor, $\sigma_{H}$ and $\sigma_{h}$, are characterized by confinement factors $M_{1}$=$\sigma_{h}$/$\sigma_{v}$=0.74 and $M_{2}$=$\sigma_{H}$/$\sigma_{v}$=0.83, with $\sigma_{h}$ orthogonal to F1, F2, and $\sigma_{H}$ orthogonal to F4 and F5. The value of $\sigma_{v}$ is computed from the formation density, assuming a hydrostatic pore pressure distribution.
@@ 4.2 Sensitivity analysis
A sensitivity analysis has been developed within the range of some geological and geomechanical quantities to investigate the configurations that are more likely to generate “unexpected” fault reactivations during the injection and storage of different kinds of fluid. The analysis addresses the role of: (i) the geological setting, by varying geometry and properties of faults and reservoir rock, (ii) the poro-elastic stress change with respect to the natural stress regime, and (iii) the space and time pore pressure gradients in the UGS formation. The analysis has been carried out in two stages:
@@
    one parameter at a time is modified with respect to the reference scenario. This enabled a thorough analysis of the individual influence of each parameter. The considered variables and scenarios are summarized in Table 2;
@@
    simultaneous variation of more than one parameter, as identified from Stage 1. The selected combination settings were tested with different injected fluids. Chemical effects were also accounted for by either weakening or hardening the Young modulus by 30% according to the type of injected fluid (Bolourinejad and Herber, 2015).
Notice that the objective of this analysis is to reproduce realistically the actual settings of the Rotliegend formation, rather then to investigate “extreme” conditions.
@@
3c Linear slip-weakening with $\varphi_{d}$ = 10 and $d_{c}=2$ mm
3d Linear slip-weakening with $\varphi_{d}$ = 20 and $d_{c}=20$ mm
@@
Table 2: Scenarios addressed in Stage 1 of the sensitivity analysis.
@@
with $\mu$ the friction coefficient, i.e., the tangent of the friction angle, $\sigma_{n}$ the normal compressive stress, and $P_{f}$ the fluid pressure in the fault. Wherever $\tau=\tau_{L}$, sliding is allowed with a potential energy release, while the condition $\tau>\tau_{L}$ is forbidden. Three parameters are used to evaluate quantitatively the fault state: (i) the criticality index $\chi=\tau/\tau_{L}$, and (ii)-(iii) the area $t_{a}$ and the sliding $d$ where $\tau=\tau_{L}$. Clearly, from (1) we have $\chi$in [0, 1]. The closer $\chi$ to 1, the more likely is a fault activation. In particular, in the sensitivity scenarios the maximum criticality factor ($\chi_{\max}$), the maximum sliding ($d_{\max}$), and the fault area ($t_{80}$) where $\chi$ exceeds 0.8 are used to lump the faults behavior.
5 Modeling Results

### v4

@@ 1.2.1 Effect on CO2
The interactions between gas, brine, and rock can lead to significant changes in the reservoir and fault properties that, in turn, can greatly impact the performance, efficiency, and safety of the storage process. However, the impact of injected gas type on the mechanical response of sedimentary rocks is still not well understood. Existing research predominantly focuses on CO2 injection (Kim and Makhnenko, 2022), with limited investigation into N2 injection (Hu et al., 2016; Fuchs et al., 2019). There is some ongoing development regarding H2 injection into sedimentary rocks (Borello et al., 2024; Vasile et al., 2024). The latest comprehensive review on H2 injection and withdrawal has been compiled by Miocic et al. (2023). From the extensive available literature on CO2-fluid-rock interactions, e.g., Rohmer et al. (2016) and Peter et al. (2022), two contrasting viewpoints emerge: CO2 can either have a negligible influence on the rock mechanical parameters, or degrade them because of mineral dissolution and modifications in pore size distribution. A stable condition is often observed when the mineral composition and pore structure remains unchanged during CO2 injection (Rimmelé et al., 2010; Bolourinejad and Herber, 2015), with the main elastic properties largely unaffected by the CO2 presence(Mikhaltsevitch et al., 2014; Hu et al., 2016). Conversely, in cases where a weakening effect is observed after CO2 injection, the mechanical strength of the rock decreases, leading to increased permeability and reduced rock stiffness (Hol et al., 2018; Manjunath et al., 2023). The dissolution of carbonate minerals, particularly calcite, reduces the cohesion of the rock matrix and weakens grain-to-grain contacts (Kim and Makhnenko, 2022). This effect has been observed in experiments on sedimentary rocks like sandstones and carbonates. Generally, carbonate-rich rocks are more susceptible to mechanical weakening due to CO2-fluid-rock interactions.
@@
While several experimental evidences support these findings, uncertainties still persist in the results. For instance, CO2-acidified brine injection can precipitate minerals, potentially increasing the strength and stiffness of the rock depending on the specific material and its mineral composition. CO2 can react with certain minerals like calcite, inducing mineral precipitation and cementation, thereby enhancing the mechanical integrity of the rock (Espinoza et al., 2018). Sandstones, for example, may experience both increase and decrease in the deformation moduli (Hangx et al., 2013; Shi et al., 2019; Tarokh et al., 2020). The variability in results regarding the effects on rock stiffness and strength highlights the need for additional experimental data to address existing uncertainties.
Regarding the geochemical effects of CO2 injection in the Upper Rotliegend formation, literature (Bolourinejad, 2015; Bolourinejad and Herber, 2015) reported an increase in permeability ranging from 10% to 30%. The increase in permeability induced by CO2 injection is often accompanied by microstructural changes, such as a change in porosity, which leads to a variation in the Young modulus, reaching up to ±30% in some cases (Harbert et al., 2020).
1.2.2 Effect on H2 and N2
@@
Regarding N2 and H2 injection, there is a limited investigation. N2 is used mainly to determine the porosity and permeability of the rock or as a control gas (Fuchs et al., 2019). Therefore, the exposure of rocks to N2 is expected to have negligible effects, although further investigations are required. In the case of H2 injection, rock degradation is expected. Hydrogen injection in UHS porous reservoirs may trigger the dissolution of carbonates and sulfates, as well as grain crushing and local compaction, possibly altering the reservoir porosity and mechanical and/or flow properties (Al-Yaseri et al., 2023). The response will be time-dependent, with pH equilibration to higher values potentially leading to further compaction and stiffening of the reservoir. However, this effect is likely to stabilize after a few cycles. Consequently, the poroelastic response of the reservoir will heavily depend on the chemical environment in the early stages of operations and become more stable and predictable in later stages. Since no data are available for Young modulus variations in the Rotliegend reservoirs, we used the reported variability for CO2 systems as a reference and assumed a ±30% change in the modeling simulations.
@@
From a geochemical point of view, gas injection may also affect faults, which represent the key factors responsible for fluid migration and containment. Gas injection can affect fault permeability, reactivation potential, and overall stability, which could lead to unintended fluid pathways or seismic activity. Concerning CO2, during the storage phase over long periods (of the order of thousands of years), fluid-rock interactions can lead to changes in fault rock fabric and mineralogy. This may result in fault frictional properties and cohesion alterations due to dissolution and cementation processes (Collettini et al., 2008). However, laboratory friction tests comparing fault rocks exposed to CO2 for several hours to samples taken from natural analogs exposed to CO2 over millions of years, show limited friction reduction due to CO2 exposure (Samuelson and Spiers, 2012). Moreover, according to the author’s knowledge, no specific laboratory studies in the Rotliegend formation show how CO2-enhanced dissolution and re-precipitation processes may impact on fault cohesion. Based on the literature, the chemical effects on fault mechanical properties after CO2 injection are considered as negligible.
@@
2  Methods and materials
We use the mathematical and numerical framework developed, discussed, and tested in Part I of this work (Franceschini et al., 2024). The modeling approach consists of one-way coupled Finite Element (FE) hydro-poromechanical simulations in fractured geological media. For ease of reference, the main modelling features relevant to the present work are briefly summarized here. The porous medium is treated as a quasi‐static, linear elastic, saturated continuum under small strains, whereas faults are represented as zero–thickness internal surfaces governed by Coulomb friction with optional slip–weakening. Normal and tangential contact tractions are enforced by Lagrange multipliers, leading to a mixed finite element formulation in which displacement is approximated with first–order hexahedral elements and fault tractions with piecewise constant interface elements. The non–linear contact problem (stick–slip–open) is solved by an active–set strategy combined with an exact Newton method and a preconditioned Krylov solver. Pore pressure fields are computed beforehand with a multiphase flow simulator and then imposed in a one–way coupled fashion, acting on the solid skeleton through Terzaghi–Biot effective stress. The same hexahedral grid is shared by the flow and geomechanical models, ensuring conservative transfer of pore pressures to the fault elements. For all the details, the reader is referred to Part I Franceschini et al. (2024).
Figure 3: (a) horizontal cross-section of the conceptual model at 2100 m depth. Faults F1 and F2 are inclined and aligned along the y-axis; F4 and F5 are vertical and aligned along the x-axis. Fault F3 separates Block 1 and Block 2, which may experience different pressure histories ($\Delta p_{1}$, $\Delta p_{2}$). (b) vertical sections of the conceptual model along the trace A-A and B-B. $\Delta\delta$ represents the possible variation in the dip angle of fault F3. This figure is not to scale to highlight the local features within the reservoir.
@@ 2.1 Conceptual reference model
The domain geometry conceptually reproduces the compartmentalization commonly observed in many UGS fields of the Rotliegend formation. The geometry and mechanical parameters were defined in agreement with the Staatstoezicht op de Mijnen (SodM) . Detailed information regarding the conceptual model is available in Franceschini et al. Franceschini et al. (2024). The reservoir compartments are approximately located at the center of a $\times30$$\times$5 km domain (Figure 3 a, Figure 4 a). The reservoir is 200-m thick, located at a depth of 2,000-m, and consists of two adjacent blocks measuring $\times2$ km each. The two reservoir compartments are separated by a fault (F3) and confined laterally by two sets of orthogonal faults (F1-F2 along the y-axis, F4-F5 along the x-axis, as shown in Figure3 a). Depending on the sealing properties of fault F3, the two compartments can have a partial hydraulic connection. The faults extend from 3,000-m to 1,600-m depth and terminate in the Zeichestein salt formation sealing the reservoir on top (Figure 2 c and Figure 3 b). Faults F1 and F2 have a dip angle of $\pm 85^{\circ}$, while faults F4 and F5 are vertical. The dip angle of fault F3 can vary from $65^{\circ}$ to $90^{\circ}$ (vertical) and from $90^{\circ}$ (vertical) to $65^{\circ}$. Block 2 can be shallower or deeper, with a maximum offset of 200 m, with respect to the Block 1.
@@
The 3D domain is discretized by 8-node hexahedra, with 253,165 nodes and 236,208 elements. A finer discretization is used between 1,800 and 2,200 m depth. Here, the reservoir element characteristic size is $\times100$$\times$20 m. The fault system embedded in the 3D grid is discretized by 5,215 zero-thickness contact elements. Standard boundary conditions with zero displacements on the outer and bottom surfaces are prescribed, whereas the top surface, representing the ground, has a stress-free condition. Figure 4 b shows a view of the discretized domain with the embedded fault system.
Figure 4: (a): axonometric view of the 3D computational grid used for the geomechanical simulation and the embedded 2D grid used to represent the fault system. Right: cut of the model along a vertical plane of symmetry with the FE mesh grid on the back part and the discretization of fault planes with IE in the front part. The various colors represent the different portions of the domain in agreement with Figure 3.
@@
The setting of the reference scenario is established on the basis of the hydro-geomechanical properties of the Rotliegend formation (Baisch et al., 2016; Buijze et al., 2017; Haug et al., 2018; Nederlandse Aardolie Maatschappij BV, 2016; Van Wees et al., 2014) and the frictional properties of faults intercepting the Rotliegend together with immediate overburden and underburden formations (Hunfeld et al., 2021, 2017, 2020). Fault F3 is kept vertical, with no offset between the reservoir blocks. In agreement with available data of the Rotliegend formation (Pijnenburg et al., 2019a; Hol et al., 2018; Zhao and Jha, 2024), the entire domain is modeled with a poroelasto-plastic constitutive law using the material properties listed in Table 3. Notice that overburden, underburden, and caprock behave elastically, as no significant stress change is computed. Faults are governed by Coulomb’s failure criterion with cohesion $c=2$ MPa and static friction angle $\varphi_{s}=30^{\circ}$. A linear slip-weakening is also possible with a reduction to the dynamic friction angle $\varphi_{d}$ at a sliding amount equal to $d_{c}$ (Franceschini et al., 2024). As to the undisturbed stress regime, the vertical effective stress, $\sigma_{v}$, is a principal component. The other two horizontal principal components of the effective stress tensor, $\sigma_{H}$ and $\sigma_{h}$, are characterized by confinement factors $M_{1}=\sigma_{h}/\sigma_{v}=0.64$ and $M_{2}=\sigma_{H}/\sigma_{v}=0.83$, with $\sigma_{h}$ orthogonal to F1, F2, and $\sigma_{H}$ orthogonal to F4 and F5. The value of $\sigma_{v}$ is computed from the formation density, assuming a hydrostatic pore-pressure distribution, and corresponds to $\sigma_{v}\approx 25$ MPa at the average reservoir depth of about 2100 m (Franceschini et al., 2024).
2.2 Fluid-dynamic model
The gas flow dynamic during injection and withdrawal is simulated by different codes according to the selected fluid: OPM Flow (The Open Porous Media Initiative, 2023; Rasmussen et al., 2021) for CH4 (UGS), and Eclipse (Schlumberger, 2014) for H2 (UHS), CO2 (CCS), and N2. For all simulations, the same well configuration is used (Staatstoezicht op de Mijnen (SodM), ). During the PP stage, two wells located next to the reservoir boundary are open in the first layer (see Fig. 5 a,b,c,d). Then, these wells are converted to injection purposes. Only for UHS, nine storage wells serve for injection and withdrawal. These wells are rather uniformly distributed throughout the reservoir, as there are no structural or petrophysical characteristics to suggest specific well locations. In Figure 5 a, the location of the injection and withdrawal wells is shown with respect to the fault system. To avoid interpolation between different computational grids, both OPM and Eclipse finite volume cells coincide with the hexahedral elements of the geomechanical model. Each reservoir block is subdivided into 4,000 regular cells with a $20\times 20\times 10$ partitioning. The horizontal and vertical permeability is $k_{h}=600$ mD and $k_{v}=300$ mD, respectively. In the flow simulations, only the reservoir layers are included in the hydraulic domain. The caprock, sideburden and underburden are not part of the flow model, and therefore their pore pressure is kept at the initial hydrostatic value throughout all simulation stages. As a consequence, all pressure variations produced during production and injection remain confined within the reservoir blocks, and the pressure field provided to the geomechanical model reflects changes occurring only inside the reservoir.
Figure 5: (a) spatial location of the injection and production wells (blue and red for injection/production for CH4, red for CO2 and N2 injection, and green for UHS cycles). (b) 3D perspective view of block 1, showing the surrounding fault planes and the subdivision into two sub-blocks, left (red) and right (blue), illustrated in the right-hand panels Right: normalized pressure distribution from the flow-dynamic simulation of CH4 (c) and H2 (d) within the right reservoir block. The grid splits vertically along the dashed line shown on the left panel to provide evidence of the pressure distribution along the vertical direction.
The PP stage is conventionally set to 10 years, with a maximum pore pressure drop of about 20 MPa. The targeted final Recovery Factor, i.e., the ratio between the cumulative produced volume of gas and the gas originally in place, is 90%, in line with theoretical reference behavior for a depletion-driven gas reservoir. Then, for CH4 the reservoir pressure recovers to the initial value in two years during the CGI phase and a set of UGS cycles, consisting of 6-month withdrawal and 6-month injection each with a 10-MPa pressure, follows. As to the other fluids, the pressure history is targeted to closely approximate the maximum static pressure drop and recovery in reservoir observed for CH4, imposing realistic storage constraints. In addition, no injection/withdrawal cycles are simulated for CO2 and N2 storage. The CO2 storage scenario considers a maximum injection rate of $4\cdot 10^{6}$ Sm3/day per well for 13 years before recovering the initial pressure value. Injection of N2 lasts 6 years with an injection rate of $4\cdot 10^{6}$ Sm3/day per well. Concerning H2, 10 years of injection/withdrawal activity are simulated after a 8-year refilling phase at a maximum rate of $4\cdot 10^{6}$ Sm3/day per well. Each UHS cycle consists of a 6-month withdrawal and a 6-month injection period, with a maximum pressure variation of 10 MPa. In all simulations, a 1-year loading step (l.s.) is used during PP and CGI/refilling, while during UGS and UHS the loading step is equal to 15 days.
It is important to recall that in the Netherlands the fluid pressure is not allowed to exceed the initial value $P_{i}$. Pore pressure propagation within the fault zone varies according to the distribution of pressure changes across different reservoir compartments. When differing pressure changes occur on opposite sides of the fault, the resulting pore pressure change within the fault, denoted as $\Delta P_{f}$, correspond to the average of the pressure changes from both sides. The fluid pressure in the fault, $P_{f}$, generates forces that oppose to the action of the pressure variation on $\sigma_{n}$.
@@ 2.3 Sensitivity analysis
A sensitivity analysis has been developed within the range of some geological and geomechanical quantities to investigate the configurations that are more likely to generate “unexpected” fault reactivations during the injection and storage of different kinds of fluid. The analysis addresses the role of: (i) the geological setting, by varying geometry and properties of faults and reservoir rock, (ii) the poro-elastic stress change with respect to the natural stress regime, and (iii) the space and time distribution of pore pressure gradients in the reservoir compartments. The analysis has been carried out in two stages:
@@
    one parameter at a time is modified with respect to the reference scenario with CH4 storage/production. This enables a thorough analysis of the individual influence of each parameter. The considered variables and scenarios are summarized in Table 4;
@@
    simultaneous variation of more than one parameter, as identified from Stage 1. To reflect potential mechanical consequences of fluid–rock chemical interactions, the Young modulus was modified by ±30% depending on the injected fluid type (Harbert et al., 2020).
Note that the objective of this analysis is to reproduce realistically the actual settings of the Rotliegend formation, rather than to investigate “extreme” conditions.
@@
3c Linear slip-weakening with $\varphi_{d}$ = 10 and $d_{c}=2$ $m$
3d Linear slip-weakening with $\varphi_{d}$ = 20 and $d_{c}=20$ $m$
@@
Table 4: Scenarios addressed in Stage 1 of the sensitivity analysis (UGS).
@@
with $\mu$ the friction coefficient, i.e., the tangent of the friction angle, $\sigma_{n}$ the normal compressive stress, and $P_{f}$ the fluid pressure in the fault. Wherever $\tau=\tau_{L}$, sliding is allowed with a potential energy release, while the condition $\tau>\tau_{L}$ is forbidden. Three parameters are used to evaluate quantitatively the fault state: (i) the criticality index $\chi=\tau/\tau_{L}$, and, where $\tau=\tau_{L}$, (ii) the reactivated area $t_{a}$, and (iii) the sliding $d$. Clearly, from Eq. (1) we have $\chi$ $\in$ [0, 1]. The closer $\chi$ to 1, the more likely is a fault activation. In particular, in the sensitivity scenarios the maximum criticality factor ($\chi_{\max}$), the maximum sliding ($d_{\max}$), and the fault area ($t_{80}$) where $\chi$ exceeds 0.8 are used to lump the faults’ behavior.
3  Analyses and results